OpenStudy (anonymous):

WILL FAN & MEDAL Given: m ∠ EGF= 60° and m∠AGF = 90°; Prove: m ∠ AGB = 30°

OpenStudy (anonymous):



OpenStudy (anonymous):

Start with what's given 1) m<EGF = 60 -given 2) m<AGF = 90 -given

OpenStudy (anonymous):

1) m<EGF = 60 -Given 2) m<AGF = 90 -Given 3) m<BGC = m<EGF -Vertical angles are congruent. 4) m<BGC = 60 -By transitive property of equality. 5) m<AGF + m<AGC = 180 -Linear pair angles are complementary. 6) 90 + m<AGC = 180 -By substitution. 7) m<AGC = 90 -By subtraction property of equality. 8) m<AGB + m<BGC = m<AGC -By angle addition postulate. 9) m<AGB + 60 = 90 -By substitution. 10) m<AGB = 30 -By subtraction property of equality.

OpenStudy (some.random.cool.kid):

i see why it says prove why Agb is 30* because I don't believe its 30*

OpenStudy (some.random.cool.kid):

agb and agf is same length. Thus being as it is they should both be a 90 degree angle.

OpenStudy (anonymous):

Yeah 😁 me too but thanks guys